OpenStudy (anonymous):

solve x+3y=0 5x+y=14

OpenStudy (anonymous):

x = -3y 5(-3y)+y=14 -15y+y=14 -14y=14 y=-1 x=-3y plug in y x=-3(-1) x=3 y=1